### v1

# ©Plug-in Authorization for Human Content Copyright Protection in Text-to-Image Model

Abstract
This paper addresses the contentious issue of copyright infringement in images generated by text-to-image models, sparking debates among AI developers, content creators, and legal entities. State-of-the-art models create high-quality content without crediting original creators, causing concern in the artistic community. To mitigate this, we propose the ©Plug-in Authorization framework, introducing three operations: addition, extraction, and combination. Addition involves training a ©plug-in for specific copyright, facilitating proper credit attribution. Extraction allows creators to reclaim copyright from infringing models, and combination enables users to merge different ©plug-ins. These operations act as permits, incentivizing fair use and providing flexibility in authorization. We present innovative approaches, “Reverse LoRA” for extraction and “EasyMerge” for seamless combination. Experiments in artist-style replication and cartoon IP recreation demonstrate ©plug-ins’ effectiveness, offering a valuable solution for human copyright protection in the age of generative AIs.
keywords:
Copyright Authorization, Copyright Protection, Text-to-image Model, Generative AI
1 Main
Figure 1: (a): ©Plug-in Authorization Process. The authorization process consists of three types of entities: user, model provider, and IP owners (artists). Users can generate copyrighted images only by accessing the relevant plug-in. The model provider offers services to users, tracks usage of plug-ins, and attributes rewards to the IP owner. The IP owners can achieve authorization by registering their ©plug-ins through addition or extraction. These ©plug-ins form a pool where users can use ©plug-ins to produce content with the IP owner’s authorization. (b): Three foundational operations achieving ©Plug-in Authorization: addition, extraction, and combination. The plug-in can be created by addition if the copyrighted work is new to the base model. Meanwhile, the plug-in can be created by extracting from the base model if the copyrighted work is already infringed by the base model. Once a pool of ©plug-ins is constructed, the combination operation can merge multiple ©plug-ins featuring the generation of multiple concepts and leave a non-infringing model excluding all multiple concepts.
Large foundation models are trained with extensive, high-quality datasets like LAION [1] and other enormous undisclosed data sources, which definitely contain copyrighted human contents. At the same time, these models not only excel at generating content based on user prompts [2, 3, 4], but also have the potential of memorizing the exact training data thanks to the huge capacity in their gigantic numbers of parameters [5].
These have sparked copyright infringement concerns among content providers, artists, and users. A notable instance is the lawsuit filed by The New York Times against OpenAI and Microsoft [6], alleging the unauthorized use of a vast number of articles for the purpose of training automated chatbots. The lawsuit seeks the destruction of the allegedly infringing chatbots and their associated training data. Similar concerns and legal actions are also emerging in the field of text-to-image generation [7].
Although it is still debatable whether using copyrighted material to train machine learning models is allowable due to the fair use exception in copyright laws [6, 7], it’s definitely that these powerful models have disrupted existing reward systems for creative artists and content providers [8]. This paper advocates revisiting the goal of copyright laws [9], which is to protect the original expression of ideas in creative works and promote scientific progress by granting authors exclusive property rights for a limited time. However, generative AI models challenge the proper attribution of rewards to copyright holders, impacting artists and domain experts who may hesitate to contribute to knowledge exchange platforms without reasonable rewards. This could lead to a shortage of fresh data for machine learning.
To address these copyright challenges in generative AI models, we introduce the “©Plug-in Authorization” framework for text-to-image generation systems (see Figure  1(a)), to align with existing Intellectual Property (IP) management practices, allowing copyright holders to register their works as plug-ins and receive rewards for their use. End users pay for generating images with copyrighted concepts, and base model providers profit from plug-in registration and usage. This framework facilitates explicit tracking of copyrighted work usage, ensuring a fair reward system.
Technically to enable an effective and efficient copyright authorization, the plug-ins, as permits, should be easily created by addition if copyrighted works are new to the base models, or by extraction if the copyrighted works are already infringed by the base model. Moreover, the plug-ins should be easily combined, which allows copyright holders to merge multiple plug-ins into a new one or enables end users to generate images with multiple copyrighted works. Meanwhile, for efficient execution, these operations should be implemented as light adaptations to the base model, e.g., parameter-efficient tuning methods or prompt designs. We introduce three foundational operations - addition, extraction, and combination - implemented using the Low-Rank Adaptor (LoRA) method [10]. These operations are essential for realizing the Copyright Plug-in Authorization (Figure 1(b) for an overview).
In summary, we make a conceptual contribution: advocating solving the problem of copyright infringement in foundation models with a ©Plug-in Authorization framework. It can offer a fair and practical solution for the attribution challenge in text-to-image generative models. We further introduce three operations addition, extraction and combination to instantiate the framework with efficient human content copyright authorizations. Technically, we propose a novel “Reverse LoRA” algorithm for extraction. It can effectively extract copyrighted concepts from the base model, achieving competitive performance for concept extraction with flexible plug-ins, which are verified with deleting various artist styles and cartoon IPs. We further propose a novel “EasyMerge” approach for combination. It is a data-free layer-wise distillation approach, which can effectively and efficiently address the challenge of combining multiple LoRA components. These methodological innovations are extensively verified with practical text-to-image generation scenarios.
2 Methods: The ©Plug-in Authorization Framework for Text-to-Image Models
As detailed in the Section 1, we implement the “©Plug-in Authorization” by utilizing the publicly available pretrained diffusion generative model, Stable Diffusion [11], along with LoRA components [10]. It is important to note that our framework is not confined to specific model structures, thereby also compatible with other foundation models such as the GPT series [12]. Additionally, it is capable of synergizing with various light fine-tuning or prompt tuning techniques [13, 14, 15, 16]. In the subsequent sections, we revisit the fundamentals of diffusion generative models and introduce the three basic operations of the framework, along with our innovative algorithms.
2.1 Preliminary on Diffusion Generative Model
Diffusion models [17, 18, 19] are probabilistic models designed to learn a data distribution. In the forward pass, Gaussian noises are successively added $T$ times to an image $X_{0}$, thereby creating a sequence ${X_{0},...,X_{T}}$ that simulates a Markov process. Conversely, the reverse process trains the model to denoise, effectively emulating the reversal of the Markov Chain. New images are generated by initially sampling random Gaussian noises and then denoising them using the model. Importantly, this process can be conditioned on inputs, such as a prompt text $c$. The denoising process, denoted as ${\Phi}_{(w)}(X_{t},c,t)$, is trained to predict the noise under the textual prompt $c$ at any timestep $t\in[0,T]$, as outlined in (1),
$\mathop{\arg\min}\limits_{w}\ \mathbb{E}_{\epsilon,X,c,t}\|{\Phi}_{(w)}(X_{t},c,t)-\epsilon\|^{2},$ (1)
where $w$ denotes the model parameters that are usually of the form U-Net [11] or DiT [20]. Recent advancements have introduced latent diffusion models as a solution to mitigate the drawbacks associated with evaluating and optimizing models in pixel space, such as low inference speed and high training costs. These latent diffusion models operate within a compressed latent space, exemplified by publicly available models such as the Stable Diffusion Model (SDM), as detailed in Rombach et al. [11]. The SDM architecture consists of a variational autoencoder (VAE) that maps images to latent space, a U-Net that learns the diffusion process, and a CLIP encoder for text embedding. Our work primarily focuses on the attention structure within the U-Net, which has been identified as the most influential component in diffusion models.
To implement the “©Plug-in Authorization”, we incorporate three foundational operations into the Stable Diffusion Model (SDM): addition, which allows copyright owners to add a plug-in for their works; extraction, which enables owners to extract a plug-in from an infringing base model; and combination, which permits users to merge plug-ins for multiple copyrighted concepts. The addition operation employs LoRA components that are added to SDM’s attention matrices, and these are then trained with copyrighted data. While the specifics of the addition operation are covered on existing model-sharing platforms like Civitai [21], our discussion will primarily focus on the extraction and combination operations in the subsequent sections.
2.2 Extraction: Reverse LoRA
Figure 2: The method of extraction consists of two steps: de-concept and re-context. The de-concept step tries to capture the target concept “Picasso” by tuning the LoRA component to match copyrighted images with the contextual prompt “The painting of building”. In the re-context step, we reverse the LoRA (so that successfully forget “Picasso”) and then further tune the LoRA with surrounding contextual prompt and non-copyrighted image pairs, to ensure the capabilities of contextual generation.
To achieve extraction, we introduce a method known as “Reverse LoRA”. This approach consists of two crucial steps to effectively extract the target copyrighted concept while maintaining the ability to generate other irrelevant concepts. The initial step, termed the de-concept step, involves removing the target concept from the base model. The subsequent step is a counterbalance to the de-concept step, involving relearning the surrounding semantic context, referred to as the re-context step. Figure 2 illustrates these two steps in the context of extracting the target concept “Picasso”.
The de-concept step. Our goal is to extract Picasso-related information from the base model to a copyright plug-in using LoRA ($w_{L}$). This involves identifying the information that represents “Picasso” via an alignment process and documenting the changes in model parameters that occur during the alignment process.
Specifically, we align copyrighted image generation, e.g., images of “the painting of a building by Picasso”, with the prompt without copyrighted text, e.g., “the painting of a building” in the base infringing model. The alignment objective is mathematically expressed as follows:
$\mathbb{E}[{\Phi}_{({w})}(\epsilon,c^{*},t)]=\mathbb{E}[{\Phi}_{({w+w_{L}})}(\epsilon,c,t)],$ (2)
where $\Phi$ is the denoising function, $w$ denotes the original network parameter, $w_{L}$ is the LoRA component, $c$ is the prompt “the painting of a building”, $c^{*}$ is the prompt “the painting of a building by Picasso”, $\epsilon$ is the initial noise and $t$ is the sampling timestep.
To achieve the alignment as defined in (2), we optimize the following objective function with respect to the LoRA parameters $w_{L}$, while keeping all other model parameters frozen. The objective function is defined as follows:
$\mathop{\arg\min}\limits_{w_{L}}\ \mathbb{E}_{\epsilon,X^{*},c,t}\|{\Phi}_{(w+w_{L})}(X^{*}_{t},c,t)-\epsilon\|^{2}$ (3)
where $X^{*}$ is the copyrighted image (or the image generated by the infringing model with the prompt “the painting of building by Picasso”), $X^{*}_{t}=\sqrt{\alpha_{t}}X^{*}+\sqrt{1-\alpha_{t}}\epsilon$ is the noisy version of $X^{*}$, $c$ is the prompt without copyright concepts, i.e., “the painting of building”, $w$ is the original network and $w_{L}$ is the LoRA weight. This optimization aims to adjust $w_{L}$ so that it effectively captures the desired information related to the copyright concept.
After adding such a LoRA component, the base model can generate Picasso-style images even when the prompts do not explicitly mention “Picasso”. Hence, the LoRA represents the Picasso style, and $w-w_{L}$ would give us a non-infringing model, which can be thought of as an analogy of a negative LoRA. However, directly using $w-w_{L}$ as the non-infringing model compromises its ability to generate images with surrounding contexts, e.g., “the painting of a building”, as shown in Figure 5 in Appendix B. This observation motivates us to further tune the LoRA with pairs of images and texts of surrounding contexts.
The re-context step. To mitigate the performance degradation of the non-infringing model when generating images with contextually related prompts, we introduce a re-context step after the de-concept step. This step involves fine-tuning the LoRA component with images and textual prompts of surrounding contexts, e.g., “the painting of a building”. To curate the dataset, we randomly generate images with the base model using the contextual prompt “the painting of a building”, while leveraging the negative prompt [22] “Picasso” to steer the generation as far away from the target concept “Picasso” as possible. Specifically, we further optimize ${w}_{L}$ with the objective,
$\mathop{\arg\min}\limits_{w_{L}}\ \mathbb{E}_{\epsilon,X,c,t}\|{\Phi}_{(w-w_{L})}(X_{t},c,t)-\epsilon\|^{2},$ (4)
[…]
Overall, after the de-concept step, the model $w-{w}_{L}$ is unable to generate images in the Picasso style, and it performs well with surrounding prompts thanks to the re-context step. Therefore, through the extraction operation, we obtain a non-infringing model $\tilde{w}=w-{w}_{L}$ and a ©plug-in ${w}_{L}$. By incorporating the ©plug-in, the model is restored to the original base model $w$, regaining the capability of successfully generating the artworks in the “Picasso” style. The intermediate results of the extraction process are visually illustrated in Figure 5 in Appendix B, showcasing the successful extraction of the targeted copyright, while preserving the model’s ability to generate images with surrounding contexts.
2.3 Combination: EasyMerge
In this section, we consider the operation combination. The combination of existing ©plug-ins becomes essential when aiming to generate an image featuring both “Snoopy” and “Mickey” concepts.
It is worth noting that simply adding these plug-ins together could yield unpredictable outcomes due to inherent correlations among these plug-ins. To facilitate the combination of multiple copyrighted concepts, we propose a novel approach named EasyMerge. This method employs a data-free, layer-wise distillation technique that only requires plug-ins and corresponding text prompts. Furthermore, with layer-wise distillation, EasyMerge achieves efficient combination in just a few iterations. The versatility of EasyMerge extends beyond the current context, potentially also applicable in other scenarios like continual learning.
Specifically, we use a new LoRA component $w_{L}$ to mimic the functionalities of each plug-in that needs to be combined. The objective is defined as follows:
$\mathop{\arg\min}\limits_{w_{L}}\ \sum_{c_{k}\in S,j\in S_{L}}\mathbb{E}_{\epsilon,t}\|{\phi^{j}_{(w+w_{L})}(\epsilon,c_{k},t)-\phi^{j}_{(w+w_{L_{k}})}(\epsilon,c_{k},t)}\|^{2},$ (5)
where $\phi^{j}$ is the output of layer $j$’s LoRA component, $w_{L_{k}}$ is the LoRA plug-in of the context $c_{k}$, ${w}_{L}$ denotes the combined plug-in, $S$ is the set of target concepts $c_{k}$ to be combined, $S_{L}$ is the set of layers that are equipped with LoRA components. Similarly to the previous section, $w$ denotes the base model parameter, $\epsilon$ is initial noise and $t$ is the sampling timestep of the diffusion process. The implementation details are described in Algorithm 1.
Algorithm 1 Combination: EasyMerge method
0:  A set $S$ of indices of plug-ins to be combined, a set $S_{L}$ of indices of layers with LoRA components, a base model $w$ and a diffusion step $T$
0:  Combined LoRA plug-in $w_{L}$
1:  while not converge do
2:     for $c_{k}\in S$ do
3:        ## The pair of plug-in and prompt to be combined, i.e., $(w_{L_{k}},c_{k})$
4:        Sample $t\sim\text{Uniform}([1...T]);\epsilon\sim N(\mathbf{0},\mathbf{I})$
5:        Record the input and output of each layer $j$: $I_{w_{L_{k}}}^{j},O_{w_{L_{k}}}^{j}\leftarrow\phi^{j}_{(w+w_{L_{k}})}(\epsilon,c_{k},t)$
6:        Compute the output of $w_{L}$: $O_{w_{L}}^{j}\leftarrow\phi_{{w_{L}}}^{j}(I_{w_{L_{k}}}^{j})$
7:        Compute the loss: $\mathcal{L}\leftarrow\sum_{j\in S_{L}}\|O_{w_{L}}^{j}-O_{w_{L_{k}}}^{j}\|^{2}$
8:        Update the parameter: $w_{L}\leftarrow w_{L}-\nabla_{w_{L}}\mathcal{L}$
9:     end for
10:  end while
We note that the above combination process can also be applied to the scenario of extracting multiple concepts simultaneously. The non-infringing model $w-w_{L}$ represents a model that simultaneously excludes multiple contexts $c_{k}$ given each extraction plugin $w_{L_{k}}$, which is called the combination of extraction.
3 Results
Given that the addition operation has already been well demonstrated by existing practices, we focus on evaluating the extraction and combination operations. We choose two typical scenarios of copyright infringement: artist-style replication and cartoon intellectual property (IP) recreation.
3.1 Experiment Setup, Metrics and Baseline Comparison
Experiment Setup. Our experiments involve fine-tuning the U-Net architecture of the Stable Diffusion Model v1.5. For extraction, we generate data with pre-trained models, using ChatGPT to create content related to both artistic styles and specific IP characters. We conduct a 10-epoch training with 30 iterations per epoch, employing a learning rate of 1.5e-4 and a rank of 40 for LoRA. The combination operation uses a learning rate of 1e-3 and a rank value of 140 for LoRA.
Evaluation Metrics. The effectiveness of extraction is measured by the discrepancy between images generated by the base model and the non-infringing model post-extraction, using the same prompts. We employ Kernel Inception Distance (KID) [23] and Learned Perceptual Image Patch Similarity (LPIPS) [24] as objective metrics, alongside human judgment, to evaluate the quality and stylistic consistency of generated images.
Baseline Comparison. We compare our extraction operation with the concept ablation approach [25] and Erased Stable Diffusion (ESD) [26], which achieve concept removal by aligning latent representations of target concepts with those of anchor concepts. Due to the complexity of image generation setups, we conservatively compare our methods with them by using similar scenarios and metrics.
3.2 Extraction and Combination of Artists’ Styles Replication
Figure 3: Results of style replication. In Figure (a), we show samples after extracting the different artist’s style in each column. Each non-infringing model exhibits a deficiency in one style generation ability, with all other style generation capabilities remaining unaffected. In Figure (b), we present samples generated after using certain ©Plug-ins in each column. Each of these ©Plug-ins serves to exclusively restore the generation of one particular style, while the generation of other styles continues to exhibit diminished performance.
Extract artist styles from the Stable Diffusion V1.5 (base model). We consider three renowned artists: (1) Vincent van Gogh, (2) Pablo Ruiz Picasso and (3) Oscar-Claude Monet. The results of individual style extraction are visually presented in Figure 3(a), where we compare images generated by the base model and the non-infringing model using target and surrounding style prompts. The images enclosed in red boxes represent the target styles, showcasing a marked contrast with the base model’s outputs. This contrast confirms the successful isolation of the target styles. Meanwhile, images for surrounding styles maintain a high degree of similarity with the base model, indicating the preservation of image quality for non-target styles.
Table 1: Quantitative comparison with baselines in artist-style extraction. Compared to Concepts-Ablation, ours extracts the target style more thoroughly, and compared to ESD, ours enjoys less damage to surrounding styles.
Metrics Methods Target style $\uparrow$ Surrounding style $\downarrow$
KID×$\times 10^{3}$ Extraction (Ours) 187 32
Concepts-Ablation 42 12
LPIPS Extraction (Ours) 0.31 0.14
ESD 0.38 0.21
Quantitative analysis, presented in Table 1, further validates the effectiveness of our extraction operation. The Kernel Inception Distance (KID) metric [23] quantifies the discrepancy between the set of images generated by the base model and that generated by the non-infringing model after extraction with the same set of prompts. KID is akin to the Fréchet Inception Distance (FID) [27] but is considered to exhibit less bias. The KID of our extraction method significantly outperforms the Concepts-Ablation method [25] by increasing from 42 to 187 for the target style, indicating more effective removal of target concepts. Additionally, our method reduces the Learned Perceptual Image Patch Similarity (LPIPS) from 0.21 to 0.14 for surrounding styles in comparison to the Erasing Stable Diffusion (ESD) method [26], indicating less degradation and a better preservation of the artistic quality. LPIPS is a robust measurement tool that effectively captures differences in human perception between two images, offering a comprehensive evaluation of stylistic variations in generated artworks.
Combination of extractions. We sequentially extract the styles of Van Gogh, Picasso, and Monet, and build a non-infringing model by using combination of extractions. Then we reintroduce them by using their corresponding copyright plug-ins. The non-infringing model, shown in the leftmost column of Figure 3(b), generates images that are notably different from those of the base model, highlighting the effectiveness of the multiple extraction operations.
Upon using the individual style copyright plug-ins, the images within the red boxes exhibit the target styles with a striking resemblance to the base model’s outputs. This outcome demonstrates that ©plug-ins can restore the model’s capability to generate artwork in the target styles while adhering to copyright restrictions.
3.3 Extraction and Combination of Cartoon IPs Recreation
In the realm of intellectual property (IP) recreation, extraction and combination operations showcase significant capabilities. Figure 4(a) presents the extraction results for three iconic IP characters: Mickey Mouse, R2D2, and Snoopy. The red-boxed images, generated after extracting the target IP prompts, deviate markedly from the base model’s outputs, demonstrating the effectiveness of target extraction. Notably, other IP images maintain similarity with the base model, confirming that the extraction is precise and does not disrupt the generation of surronding IP-related content.
Quantitative analysis in Table 2 reveals a substantial 2.6-fold increase in the Kernel Inception Distance (KID) score for the targeted IPs, while other IPs’ KID metrics remain stable, underscoring the method’s targeted isolation of specified IPs.
Figure 4: Results of IP Recreation. In Figure (a), each column on the right represents the output after extraction. We successfully extract the unique IPs of Mickey, R2D2, and Vader independently, preserving the generation of other IPs. In Figure(b), we generate either Mickey or Vader in a single image by using the corresponding Plug-in and generate them both by using the combined ©plug-in.
Moreover, our extensive large-scale experiments confirm that the extraction process does not affect the model’s capacity to produce content unrelated to particular IPs. The model’s performance post-extraction remains stable across routine or common usage scenarios, as elaborated in the Appendix C. This guarantees that while the model adheres to copyright limitations by appropriately handling or isolating particular IPs, it maintains its overall usefulness and diverse creative capabilities.
Table 2: Quantitative comparison in IP recreation. We increase the KID of the target IP about 2.6 times compared with Concepts-Ablation while keeping the KID of the surrounding IP on par.
Metrics Methods Target IP $\uparrow$ Surrounding IP $\downarrow$
KID×$\times 10^{3}$ Extraction (Ours) 131 17
Concepts-Ablation 50 15
The combination operation is illustrated in Figure 4(b), where the initial image lacks recognizable IPs. The subsequent images, using individual and combined copyright plug-ins for Mickey and Vader, clearly display the respective IPs. The final image is generated upon using the combined ©plug-in, successfully displaying both IPs. This sequence highlights the plug-in’s efficacy in selectively and collectively restoring the model’s ability to generate IP-specific content.
These results indicate the efficacy of combining multiple extractions, where the non-infringing model’s ability to produce images themed with either Mickey Mouse or Darth Vader is disabled. With the use of the respective copyright plug-in, the model regains the capability to create content related to the specific IP.
4 Discussion
The rapid advancement of generative AI models has sparked significant concern over their potential to produce copyright-infringing content. As these models elevate the quality of their outputs, the lack of proper attribution to human creators becomes increasingly problematic. To mitigate these societal concerns, we introduce the “Copyright Plug-in Authorization” framework, a system designed to align with the principles of copyright law and ensure that the use of copyrighted material in AI-generated content is appropriately managed and rewarded.
One instantiation of our framework leverages the Low-Rank Adaptor (LoRA) to convert copyrighted data into plug-ins, facilitating the tracking of content usage and the fair distribution of rewards to original creators. This approach not only respects the rights of content creators but also provides a clear and accessible mechanism for users to engage with copyrighted material in a legally compliant manner.
Nonetheless, the current framework may still face challenges when managing a vast array of plug-ins. As the number of plug-ins grows, maintaining a user-friendly interface and ensuring seamless access to specific generations becomes critical. This requires innovative solutions to organize and retrieve plug-ins effectively, without compromising the user experience. Another challenge is ensuring the backward compatibility of plug-ins when the base model is updated. Retraining the entire collection of plug-ins following a model update can be prohibitively expensive and time-consuming. Ensuring backward compatibility, so that previously created plug-ins continue to function with new model iterations, is essential to avoid obsolescence and maintain user trust.
Our Copyright Plug-in Authorization framework represents an initial step towards balancing the innovative potential of generative AI with the imperative of copyright protection. By addressing the challenges of plug-in management, model updating, and performance maintenance, we aim to foster an ecosystem where creativity and legal compliance can coexist and thrive.
Appendix A Related Work
In order to position our work in the vast literature, we review related work through two perspectives: scope and technique. It is worthy to note that some of the literature touch both sides and we organize them in a way most related to ours.
A.1 Scope Related: Copyright, Data Contribution and Credit Attribution
Recent text-to-image generative models are trained with large scale datasets [1, 28], which cannot be guaranteed free of copyrighted data. At the same time, the state-of-the-art models are capable of generating high-quality and valuable creative images comparable to human creators or even memorizing the data points in the training set [29], which arouses copyright concerns about the training data and brings anxiety to the artist community.
[…]
Shan et al. [35] propose Image Cloaking that suggests adding adversarial perturbations before posting artistic works on the internet so as to make them unlearnable for machine learning model, which has been pointed out to be hard to defend against future learning algorithms [36].
Theoretically, [37, 38] connect the copyright protection of training data with the concept of differential privacy and discuss their subtle differences. [39] further formulate the copyright problem with a near free access (NAF) notion to bound the distance of the generative distributions of the models trained with and without the copyrighted data.
Our paper distinguishes largely from all previous works as we do not try to prohibit generating copyrighted concepts but instead we introduce a copyright authorization for the generative model to reward the copyright owners with fairness and transparency. From this aspect, our paper is also related with literature of monetizing the training data [40, 41, 42, 43] or attributing credits for the generative contents [8], but we establish a very distinct way to reward the authorship.
Heated discussion is also around the copyright for AI generated art work Franceschelli and Musolesi [44], Abbott and Rothman [45]. The Review Board of the United States Copyright Office has recently refused the copyright registration of a two-dimensional AI generated artwork entitled “A Recent Entrance to Paradise”. However, [45] argues for giving the copyright to AI generated works, which will encourage people to develop and use creative AI, promote transparency and eventually benefit the public interest.
A.2 Technique Related: Concept Removal, and Negative Sampling
Our extraction operation is closely related with the concept removal for generative models. [26, 25] remove target concepts by matching the generation distribution of contexts with target concepts and that of contexts without target concepts. [46] forget target concepts by minimizing the cross attention of target concepts with that of target images. Heng and Soh [47] leverage the reverse process of continual learning to promote the controllable forgetting of target contents in deep generative models.
We note that negative sampling [22] can also prevent generating certain concepts. Specifically, end users can set conditional context and negative context to guide the diffusion process to generate images conforming the conditional context while being far away from the negative context. Only negative sampling cannot stop copyright infringing generation because the contexts are set freely and adversarially by end users.
In contrast, for a specific copyrighted concept, our extraction operation takes an “Reverse LoRA” approach to disentangle the base model into two parts: the non-infringing base model and the plug-in LoRA component for the copyrighted concept. Specifically, we use negative sampling to generate non-infringing images, which serve as training data for copyright plug-in. From the aspect of parameter efficient fine-tuning, our paper is related to literature [48, 49, 50, 10, 51].
Our combination operation is related with the widely studied “knowledge distillation” [52, 53, 54, 55, 56], but entails large difference from previous work. We combine multiple copyright plug-ins that are LoRA components for different targets, and we take data free approach due to practical constraint.
Appendix B Intermediate results of extraction
We present an intricate analysis of the extraction process, elucidating distinct phases that delineate the evolutionary trajectory of the non-infringing model’s performance. Following the de-concept step, the model experiences a transient phase marked by a temporary impairment in its ability to generate semantically meaningful images. Subsequently, the re-context step engenders a noticeable restoration, enhancing the model’s proficiency in generating semantically rich images. Importantly, despite this recuperation, the model retains its inherent limitation – the incapacity to generate artwork in the distinctive style synonymous with Picasso. This observation underscores the success of the de-concept step, wherein the LoRA component effectively captures and removes the target concept, leading to the temporary impairment in the non-infringing model. The subsequent re-context step rectifies this performance decrease without reintroducing any information about the target concept. In essence, the extraction process successfully achieves the targeted concept extraction.
Figure 5: Intermediate results of extraction. After the de-concept step, the non-infringing model’s generative abilities become significantly limited, predominantly manifesting as the production of noise. After the re-context step, the generation’s prowess is rejuvenated, but due to the absence of learning Picasso-style images, the model remains unable to generate artwork in the style of Picasso.
Appendix C Experiment on ordinary objects generation
We evaluated the influence of extraction on the generation of ordinary objects. We utilize 5000 textual captions selected from the validation set in MS-COCO [57] as prompts, generating 5000 images using SD1.5 and the non-infringing model that extracts R2D2 and Picasso, respectively. Several randomly selected images are displayed in Figure 6. For illustrative purposes, we also generated results for concept-ablation and ESD on MS-COCO, respectively. Images within the same column exhibit substantial similarity, indicating that extraction does not exert an impact on the generation of ordinary items.
Table 3: Quantitative results on MS-COCO. FID and KID metrics for removing the Picasso style are presented in the upper two rows, while those for removing R2D2 are displayed in the lower two rows.
Domain Method FID $\downarrow$ KID×$\times 10^{3}$$\downarrow$
Style replication Extract Picasso 24.04 2.83
Erase Picasso 25.20 3.39
IP recreation Extract R2D2 20.55 2.36
Ablate R2D2 18.97 1.34
Figure 6: Ordinary objects generation after extraction. Row 1 displays images generated by Stable Diffusion V-1.5. Rows 2 and 3 illustrate images generated after the removal of the IP character R2D2, while Rows 4 and 5 showcase images generated after the elimination of Picasso’s style. Rows 3 and 5 serve as the baseline, representing concept-ablation and ESD, respectively. Notably, after the extraction of R2D2 and Picasso, the non-infringing model retains the capability to generate commonplace objects sourced from the MS-COCO dataset [57].
As depicted in Table3, we calculate some quantitative metrics like FID and KID. It is noteworthy that Concept-Ablation [25] only releases the checkpoint of ablating “R2D2” and ESD [26] only releases the checkpoint of erasing “Picasso”. Thus, we compare them separately by using their respective checkpoints. Evaluation metrics consistently maintain low values, further affirming that extraction does not compromise the generation of ordinary items.